The surface area of a cube is 294 cm squared. What is the length of its side?
Answer by mananth(16949)

the cube has 6 faces
let length of side be l
area of one face = l^2
area of 6 faces = 6l^2
6l^2=294
l^2=294/6
l^2=49
l=7 cm
..
m.ananth@hotmail.ca